if sinA+sinB = root3 (cos B-cos A) them sin3B+sin3A =

if sinA+sinB= root3(cos B-cos A) them sin3B+sin3A =

Grade:11

1 Answers

Askiitians_Expert Yagyadutt
13 years ago

Hello Hemanth

 

SinA + sinB = root3(cosB) - root3(cosA)

 

sinA + root3.cosA  =  root3cosB - sinB

 

Multiply by 1/2

 

1/2 sinA + root3/2 cosA = root(3)/2 cosB - 1/2 sinB

 

sin ( A + 60) = sin ( 60 - B)

 

so A + 60 = 60 - B

 

or  A = -B

 

So  Sin3B + sin3A = sin3B - sin3B = 0

 

Ans

 

With regards

Yagya

askiitians_expert